OpenStudy (anonymous):

Multiply by the method of your choice. (y - 6)^3

OpenStudy (anonymous):

@SithsAndGiggles Can you help me with a question

OpenStudy (anonymous):

\[\left( a-b \right)^3=a^3-b^3-3ab(a-b)\]

OpenStudy (anonymous):

How do plug the problem into the equation

OpenStudy (anonymous):

I believe I have to use this formula: (A+B)^2